Given that U

 is a function of x,y

 and z

 and A

 a vector field, prove that:


∇×(UA)=(∇U)×A+U(∇×A).



1
Expert's answer
2021-12-15T09:12:21-0500

Given:- U is a scalar function of x, y, z and A is a vector field

Let "A_{1}, A_{2}, A_{3}" are components of A

 "\therefore \nabla \cdot(U A)=\frac{\partial}{\partial x}\left(U A_{1}\right)+\frac{\partial}{\partial y}\left(U A_{2}\right)+\frac{\partial}{\partial z}\left(U A_{3}\right)"  

using product rule

"\begin{aligned}

\nabla \cdot(U A) &=\left[\left(\frac{\partial U}{\partial x}\right) A_{1}+U \frac{\partial A_{1}}{\partial x}\right]+\left[\left(\frac{\partial U}{\partial y}\right) A_{2}+U\left(\frac{\partial A_{2}}{\partial y}\right)\right] \\

&+\left[\left(\frac{\partial U}{\partial z}\right) A_{3}+U\left(\frac{\partial A_{3}}{\partial z}\right)\right] \\

&=\frac{\partial U}{\partial x} A_{1}+\frac{\partial U}{\partial y} A_{2}+\frac{\partial U}{\partial z} A_{3}+U\left(\frac{\partial A_{1}}{\partial x}+\frac{\partial A_{2}}{\partial y}+\frac{\partial A_{3}}{\partial z}\right) \\

&=\left(\frac{\partial U}{\partial x} \hat{\imath}+\frac{\partial U}{\partial y} \hat{j}+\frac{\partial U}{\partial z} \hat{k}\right) \cdot\left(A_{1} \hat{i}+A_{2} \hat{\jmath}+A_{3} \hat{k}\right)+U(\nabla \cdot A)

\end{aligned}"

we recognize

"\nabla \cdot(\cup A)=(\nabla \cup) \cdot A+U(\nabla \cdot A) \\

\text { Hence proved }"